# Regular Vaidya solutions of effective gravitational theories

### v1

@@ I Introduction
It is widely expected that a full solution of these difficulties requires unifying quantum field theory and general relativity. While existing theories of quantum gravity are still far from providing a complete picture Carballo-Rubio:2025fnc , the leading low-energy effects of such theories can often be described in terms of effective geometric modifications of general relativity (recent reviews of gravitational collapse in quantum gravity include Eichhorn:2022bgu ; Ashtekar:2023cod ; Platania:2023srt ). Research along this line has, however, stumbled upon substantial difficulties. There is evidence that regularizing general black hole interiors requires field equations of order higher than second Knorr:2022kqp ; Borissova:2023kzq , but higher-order field equations pose important challenges even for basic working notions such as well posedness Kovacs:2020pns ; Kovacs:2020ywu ; Brady:2023dgu .
@@
In general relativity, the Vaidya solution is a generalization of the Schwarzschild solution with null dust (radiation) as a matter source which modifies the mass of the central black hole Vaidya:1951fdr ; Vaidya:1953zza ; Vaidya:1966zza , or its time-reversal in which the central object is instead a white hole. Due to its simplicity and physical significance, this key solution is used extensively in discussions of black hole physics across different levels, from textbooks and lecture notes on the subject Strominger:1994tn ; Poisson:2009pwt to research papers discussing core aspects of black hole physics, including diverse definitions of black hole boundaries Schnetter:2005ea ; Ben-Dov:2006zmw ; Bengtsson:2008jr ; Nielsen:2010wq and the phenomena of mass inflation Ori:1991zz ; Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca and black hole evaporation Parikh:1998ux ; Hayward:2005gi ; Frolov:2014jva . Showing that an equally simple generalization of this famous solution exists for second-order effective gravitational theories, including cases describing dynamical regular black holes instead, is a crucial development in the study of physics beyond general relativity.
@@ II Second-order effective gravitational actions
The dynamics of the $D$–dimensional metric $g^{(D)}_{\mu\nu}$ and any other matter fields is described by a $D$–dimensional action $\mathscr{S}^{(D)}=\int\text{d}^{D}y\sqrt{-g^{(D)}}\left[\mathscr{L}_{\rm G}^{(D)}/\kappa+\mathscr{L}_{\rm M}^{(D)}\right]$, where $\mathscr{L}_{\rm G}^{(D)}$ and $\mathscr{L}_{\rm M}^{(D)}$ are the gravitational and matter Lagrangian densities, respectively, and $\kappa=16\pi$ in natural units. The equations of motion can be obtained by a variational principle and then restricted to spherically symmetric metrics of the form in Eq. (1). However, due to the compact nature of the rotation group, it is equivalent to restrict the $D$–dimensional actions to spherically symmetric metrics of the form in Eq. (1) and then perform variations Palais:1979rca ; Fels:2001rv ; Frausto:2024egp . This is the route we will follow here.
When restricting the action $\mathscr{S}^{(D)}$ to spherically symmetric metrics, the resulting gravitational Lagrangian is $\Omega_{(D-2)}\mathscr{L}_{\rm G}$, where $\Omega_{(D-2)}=\int\text{d}\Omega_{(D-2)}=2\pi^{(D-1)/2}/\Gamma(D/2-1/2)$ is a numerical constant and $\mathscr{L}_{\rm G}$ a function of a 2–dimensional metric $g_{ab}(x)$ and a scalar field $R(x)$, corresponding to general 2–dimensional dilaton (2DD) gravities Colleaux:2019ckh . In general, the equations of motion will be of arbitrarily high order, depending on the specific $D$–dimensional theory considered. This space of theories is completely general and includes the dynamics of the spherically symmetric sector of any $D$–dimensional metric theory that admits a variational principle formulation.
@@
Expressions for the variations of the Lagrangian density in Eq. (2) are known Kunstatter:2015vxa ; Colleaux:2019ckh . Adding matter, we can write down the following three independent field equations:
@@
as well as the conserved $2$–dimensional source
@@
with the $D$–dimensional stress-energy tensor $T^{(D)}_{\mu\nu}$ defined using the standard Hilbert prescription.
Solutions of the field equations (4) in vacuum ($T_{ab}=0$) are known (e.g., Ziprick:2010vb ; Kunstatter:2015vxa ). However, to the best of our knowledge, it is the first time that exact solutions in the presence of a source, describing dynamical black holes, are studied in this formalism.
@@ III Vaidya solutions
The analysis above includes any metric gravitational theory with second-order field equations in spherical symmetry, recovering known results for theories such as Lovelock gravity Kobayashi:2005ch ; Maeda:2005ci ; Dominguez:2005rt ; Ghosh:2008jca ; Cai:2008mh but also proving the existence of previously unknown Vaidya solutions in other frameworks such as quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which will be discussed in the examples below. It can also accommodate Vaidya geometries obtained in studies of effective quantum gravity corrections, regardless of whether these are singular or regular Borissova:2022mgd ; Delaporte:2024and . While in the following we focus on regular geometries, this wide range of potential applications is interesting in its own right.
@@ IV Regular Vaidya solutions
The on-shell function $\Theta(R,f)$ does not depend on $R$, as shown in Eq. (13), and it is equivalent to the mass function defined in Kunstatter:2015vxa .
@@ V Examples
The first example are Ziprick-Kunstatter gravities, defined as the $D$–dimensional generalization of the 4-dimensional theories studied by Ziprick and Kunstatter Ziprick:2010vb (see also Kunstatter:2015vxa ; Barenboim:2024dko ; Barenboim:2025ckx ), with $H_{2}=\text{d}^{2}\Phi(R)/\text{d}R^{2}$, $H_{3}=\text{d}\Phi(R)/\text{d}R$ and $H_{4}=\Phi(R)$, with a a function $\Phi(R)$ behaving as $\Phi(R)\simeq R^{D-2}$ for $R\rightarrow\infty$ [thus recovering general relativity at large distances]. We thus have $\alpha=\text{d}^{2}\Phi/\text{d}R^{2}(1-X)$, $\beta=-\text{d}\Phi/dR$ which, in accordance with the regularity conditions discussed above, implies that $\Phi(R)$ must diverge at least as $\Phi(R)\propto R^{-1}$ for $R\rightarrow 0$. We also have $\Theta(R,X)=(1-X)\text{d}\Phi/\text{d}R$ and $f(V,R)=1-2(D-2)M(V)/\left(\text{d}\Phi/\text{d}R\right)$.
The second example are quasitopological gravities, higher-curvature theories defined for $D\geq 5$ and displaying second-order field equations on spherically symmetric backgrounds Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl , which are thus included in our analysis. To show this explicitly, we note that the two cases above are particular instances of a general expression $\Theta(R,X)=(D-2)R^{D-1}\Psi(Z)$, where $\Psi(Z)$ is a function of the variable $Z=(1-X)/R^{2}$. We then obtain $\alpha=(D-2)R^{D-2}\left[(D-1)\Psi-2Z\text{d}\Psi/\text{d}Z\right]$ and $\beta=-(D-2)R^{D-3}\text{d}\Psi/\text{d}Z$. These expressions coincide with the ones obtained when considering the spherically symmetric reduction of quasitopological gravities Bueno:2024eig ; Bueno:2024zsx ; Bueno:2025gjg .
@@ VI Singularity regularization and energy transfer
The matter source in Eq. (8) is singular at $R=0$. However, the modified Vaidya solutions we consider here are explicitly regular by construction, whenever the conditions in Eq. (14) are satisfied.
@@
The only component of the matter stress-energy tensor $T^{(D)}_{\mu\nu}$ is an ingoing radial null flux $T^{(D)}_{VV}$ that is divergent at $R=0$ and has a constant integral over all spheres throughout its collapse, $\Omega_{(D-2)}T_{VV}=\Omega_{(D-2)}R^{D-2}T_{VV}^{(D)}$. On the other hand, the $\mathcal{T}^{(D)}_{VV}$ component of the effective stress-energy tensor contains an ingoing flux proportional to $\dot{M}$ that is also divergent at $R=0$, but with an overall integral over spheres that depends on the radius. Indeed, the piece proportional to $\dot{M}$ in $\Omega_{(D-2)}\mathcal{T}_{VV}=\Omega_{(D-2)}R^{D-2}\mathcal{T}_{VV}^{(D)}$ vanishes in the limit $R\rightarrow\infty$ and is switched on progressively as the radius decreases, until becoming $-\Omega_{(D-2)}T_{VV}$ exactly at $R=0$. Due to $\mathcal{T}^{(D)}_{\mu\nu}$ being conserved, this ingoing flux of energy must be created at the expense of leaving behind a residual distribution of energy, encoded in the remaining non-zero pieces of $\mathcal{T}^{(D)}_{\mu\nu}$, which describe how the energy in the collapsing or accreting null dust is stored in the gravitational field of the regular black hole after being effectively transferred.
@@
Such geometric constructions have frequently been used to model the expected effects of semiclassical backreaction, and are ubiquitous in the discussions of the evaporation of regular black holes, including Hayward:2005gi ; Frolov:2014jva ; Frolov:2017rjz ; Cardoso:2023guh ; Carballo-Rubio:2024dca . Here these arise as exact solutions, which opens the possibility of a thorough understanding of the dynamical properties of evaporating regular black holes, such as their behavior under radial perturbations and the backreaction of the mass inflation instability Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca . Adding semiclassical sources, along the lines of the recent analysis for Reissner–Nordström black holes Boyanov:2025otp (see also Barcelo:2020mjw ; Barcelo:2022gii ), would also allow for a self-consistent modeling of the semiclassical dynamics of these regular objects.
Figure 1: Left: Vaidya solution in general relativity representing the formation of a Schwarzschild black hole from the collapse of null dust originating from past null infinity, as encoded in the flux component of $T_{\mu\nu}^{(D)}$, represented with blue arrows and contained within the blue dashed lines. The trapped region, characterized by $f(V,R)<0$, is shaded in gray. Right: regular Vaidya solution representing the formation and disappearance of a regular black hole due to the collapse and accretion of null dust. Aside from the classical matter source $T^{(D)}_{\mu\nu}$, the effective source $\mathcal{T}^{(D)}_{\mu\nu}$ has two distinct components: an ingoing flux (red arrows) with energy of opposite sign to that of the null dust, and the energy distribution deposited in the regular black hole (red shaded region). Both classical matter and effective sources cancel identically at $R=0$. The energy stored in the regular black hole grows while $\dot{M}>0$ (in between the blue dashed lines), decreases while $\dot{M}<0$ (in between the green dashed lines), and remains constant otherwise.
@@ Acknowledgements.
The authors are grateful to Jose Beltrán Jiménez for insightful discussions on 2–dimensional Horndeski gravities, and to Carlos Barceló, Vitor Cardoso and Stefano Liberati for feedback on the first version of the manuscript. RCR acknowledges financial support provided by the Spanish Government through the Ramón y Cajal program (contract RYC2023-045894-I), the Grant No. PID2023-149018NB-C43 funded by MCIN/AEI/10.13039/501100011033, and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033, as well as the hospitality of the Center of Gravity, a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. VB acknowledges support from the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587, as well as from the Spanish Government through the Grants No. PID2020-118159GB-C43, PID2020-118159GB-C44, PID2023-149018NB-C43 and PID2023-149018NB-C44 (funded by MCIN/AEI/10.13039/501100011033).

### v2

@@ I Introduction
It is widely expected that a full solution of these difficulties requires unifying quantum field theory and general relativity. While existing theories of quantum gravity are still far from providing a complete picture Carballo-Rubio:2025fnc , the leading low-energy effects of such theories can often be described in terms of effective geometric modifications of general relativity (recent reviews of gravitational collapse in quantum gravity include Eichhorn:2022bgu ; Ashtekar:2023cod ; Platania:2023srt ). Research along this line has, however, stumbled upon substantial difficulties. There is evidence that regularizing general black hole interiors requires field equations of order higher than second Knorr:2022kqp ; Borissova:2023kzq , but higher-order field equations pose important challenges even for basic working notions such as well-posedness Kovacs:2020pns ; Kovacs:2020ywu ; Brady:2023dgu .
@@
In general relativity, the Vaidya solution is a generalization of the Schwarzschild solution with null dust (radiation) as a matter source which modifies the mass of the central black hole Vaidya:1951fdr ; Vaidya:1953zza ; Vaidya:1966zza , or its time-reversal in which the central object is instead a white hole. Due to its simplicity and physical significance, this key solution is used extensively in discussions of black hole physics across different levels, from textbooks and lecture notes on the subject Strominger:1994tn ; Poisson:2009pwt to research papers discussing core aspects of black hole physics, including diverse definitions of black hole boundaries Schnetter:2005ea ; Ben-Dov:2006zmw ; Bengtsson:2008jr ; Nielsen:2010wq , the phenomenon of mass inflation Ori:1991zz ; Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca , and black hole evaporation Parikh:1998ux ; Hayward:2005gi ; Frolov:2014jva . Showing that an equally simple generalization of this famous solution exists for second-order effective gravitational theories, including cases describing dynamical regular black holes instead, is a crucial development in the study of physics beyond general relativity.
@@ II Second-order effective gravitational actions
We start considering as motivation the dynamics of the $D$–dimensional metric $g^{(D)}_{\mu\nu}$ and any other matter fields, described by a $D$–dimensional action $\mathscr{S}^{(D)}=\int\text{d}^{D}y\sqrt{-g^{(D)}}\left[\mathscr{L}_{\rm G}^{(D)}/\kappa+\mathscr{L}_{\rm M}^{(D)}\right]$, where $\mathscr{L}_{\rm G}^{(D)}$ and $\mathscr{L}_{\rm M}^{(D)}$ are the gravitational and matter Lagrangian densities, respectively, and $\kappa=16\pi$ in natural units. The equations of motion can be obtained by a variational principle and then restricted to spherically symmetric metrics of the form in Eq. (1). However, due to the compact nature of the rotation group, it is equivalent to restrict the $D$–dimensional actions to spherically symmetric metrics of the form in Eq. (1) and then perform variations Palais:1979rca ; Fels:2001rv ; Frausto:2024egp .
When restricting the action $\mathscr{S}^{(D)}$ to spherically symmetric metrics, the resulting gravitational Lagrangian is $\Omega_{(D-2)}\mathscr{L}_{\rm G}$, where $\Omega_{(D-2)}=\int\text{d}\Omega_{(D-2)}=2\pi^{(D-1)/2}/\Gamma(D/2-1/2)$ is a numerical constant and $\mathscr{L}_{\rm G}$ a function of a 2–dimensional metric $g_{ab}(x)$ and a scalar field $R(x)$, corresponding to general 2–dimensional dilaton (2DD) gravities (where using the term dilaton stresses its geometric origin). In general, the equations of motion will be of arbitrarily high order, depending on the specific $D$–dimensional theory considered. This space of theories is completely general and includes the dynamics of the spherically symmetric sector of any $D$–dimensional metric theory that admits a variational principle formulation.
@@
Expressions for the variations of the Lagrangian density in Eq. (2) are known (see the detailed discussion in Colleaux:2019ckh ). For our purposes here, it suffices to write explicitly the following three independent field equations in the presence of matter:
@@
as well as the $2$–dimensional source
@@
where the $D$–dimensional stress-energy tensor $T^{(D)}_{\mu\nu}$ is defined using the standard Hilbert prescription.
@@ III Vaidya solutions
When considering the particular cases in which $M(V)$ becomes a constant, we recover the vacuum solutions in Kunstatter:2015vxa . The most general Lagrangian density introduced in Kunstatter:2015vxa differs from Eq. (2) by a total derivative (see Colleaux:2019ckh ), which in particular implies that the theories defined by Eq. (2) must also satisfy Birkhoff’s theorem.
@@
The analysis above includes any metric gravitational theory with second-order field equations in spherical symmetry, recovering known results for theories such as Lovelock gravity Kobayashi:2005ch ; Maeda:2005ci ; Dominguez:2005rt ; Ghosh:2008jca ; Cai:2008mh but also proving the existence of previously unknown Vaidya solutions in other frameworks such as quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl ; Bueno:2024dgm , which will be discussed in the examples below. It can also accommodate Vaidya geometries obtained in studies of effective quantum gravity corrections, regardless of whether these are singular or regular Borissova:2022mgd ; Delaporte:2024and . While in the following we focus on regular geometries, this wide range of potential applications is interesting in its own right.
@@ IV Regular Vaidya solutions
The on-shell function $\Theta(R,f)$ does not depend on $R$, as shown in Eq. (13).
@@ V Examples
The first example are the $D$–dimensional generalization of the theories defined for 4-dimensional spherically symmetric spacetimes by Ziprick and Kunstatter Ziprick:2010vb (see also Kunstatter:2015vxa ; Barenboim:2024dko ; Barenboim:2025ckx ), namely $H_{2}=\text{d}^{2}\Phi(R)/\text{d}R^{2}$, $H_{3}=\text{d}\Phi(R)/\text{d}R$ and $H_{4}=\Phi(R)$, with a a function $\Phi(R)$ behaving as $\Phi(R)\simeq R^{D-2}$ for $R\rightarrow\infty$ [thus recovering general relativity at large distances]. We thus have $\alpha=\text{d}^{2}\Phi/\text{d}R^{2}(1-X)$, $\beta=-\text{d}\Phi/dR$ which, in accordance with the regularity conditions discussed above, implies that $\Phi(R)$ must diverge at least as $\Phi(R)\propto R^{-1}$ for $R\rightarrow 0$. We also have $\Theta(R,X)=(1-X)\text{d}\Phi/\text{d}R$ and $f(V,R)=1-2(D-2)M(V)/\left(\text{d}\Phi/\text{d}R\right)$.
The second example are quasitopological gravities Myers:2010ru ; Oliva:2010eb ; Bueno:2016xff ; Hennigar:2017ego ; Ahmed:2017jod ; Bueno:2019ycr ; Bueno:2022res ; Moreno:2023rfl ; Bueno:2024dgm , higher-curvature theories defined for $D\geq 5$ and displaying second-order field equations for spherically symmetric spacetimes, which are thus included in our analysis. To show this explicitly, we note that the two cases above are particular instances of a general expression $\Theta(R,X)=\Psi(Z)R^{2}\text{d}\Phi/\text{d}R$, where $\Psi(Z)$ is a function of the variable $Z=(1-X)/R^{2}$. For $\Phi(R)=R^{D-2}$, we obtain $\alpha=(D-2)R^{D-2}\left[(D-1)\Psi-2Z\text{d}\Psi/\text{d}Z\right]$ and $\beta=-(D-2)R^{D-3}\text{d}\Psi/\text{d}Z$. These expressions coincide with the ones obtained when considering the spherically symmetric reduction of quasitopological gravities for the study of the gravitational collapse of timelike thin shells Bueno:2024eig ; Bueno:2024zsx and timelike dust Bueno:2025gjg in $D\geq 5$ (embedding specific regular black holes in quasitopological gravities involves additional constraints that prevent specific metrics to be realized in some dimensionalities).
@@ VI Singularity regularization and energy transfer
The matter source in Eq. (8) is singular at $R=0$. However, the corresponding Vaidya solutions are explicitly regular whenever the conditions in Eq. (14) are satisfied.
@@
The only component of the matter stress-energy tensor $T^{(D)}_{\mu\nu}$ is an ingoing radial null flux $T^{(D)}_{VV}$ that is divergent at $R=0$, and its integral $\Omega_{(D-2)}T_{VV}=\Omega_{(D-2)}R^{D-2}T_{VV}^{(D)}$ is constant over all spheres throughout its collapse. On the other hand, the $\mathcal{T}^{(D)}_{VV}$ component of the effective stress-energy tensor contains an ingoing flux proportional to $\dot{M}$ that is also divergent at $R=0$, but with an overall integral over spheres that depends on the radius. Indeed, the piece proportional to $\dot{M}$ in $\Omega_{(D-2)}\mathcal{T}_{VV}=\Omega_{(D-2)}R^{D-2}\mathcal{T}_{VV}^{(D)}$ vanishes in the limit $R\rightarrow\infty$ and is switched on progressively as the radius decreases, until becoming $-\Omega_{(D-2)}T_{VV}$ exactly at $R=0$. Due to $\mathcal{T}^{(D)}_{\mu\nu}$ being conserved, this ingoing flux of energy must be created at the expense of leaving behind a residual distribution of energy, encoded in the remaining non-zero pieces of $\mathcal{T}^{(D)}_{\mu\nu}$, which describe how the energy in the collapsing or accreting null dust is stored in the gravitational field of the regular black hole after being effectively transferred.
@@
Such geometric constructions have frequently been used to model the expected effects of semiclassical backreaction, and are ubiquitous in the discussions of the evaporation of regular black holes, including Hayward:2005gi ; Frolov:2014jva ; Frolov:2017rjz ; Cardoso:2023guh ; Carballo-Rubio:2024dca . Here these can be constructed as exact solutions, which opens the possibility of a thorough understanding of the dynamical properties of evaporating regular black holes, such as their behavior under radial perturbations and the backreaction of the mass inflation instability Frolov:2017rjz ; Bonanno:2020fgp ; Carballo-Rubio:2021bpr ; Carballo-Rubio:2024dca . Adding semiclassical sources, along the lines of the recent analysis for Reissner–Nordström black holes Boyanov:2025otp (see also Barcelo:2020mjw ; Barcelo:2022gii ), would also allow for a self-consistent modeling of the semiclassical dynamics of these regular objects. Recent related numerical breakthroughs can be found in Barenboim:2024dko ; Barenboim:2025ckx .
Figure 1: Left: Vaidya solution in general relativity representing the formation of a Schwarzschild black hole from the collapse of null dust originating from past null infinity, as encoded in the flux component of $T_{\mu\nu}^{(D)}$, indicated by blue arrows and contained within the blue dashed lines. The trapped region, characterized by $f(V,R)<0$, is shaded in gray. Right: regular Vaidya solution representing the formation and disappearance of a regular black hole due to the collapse and accretion of null dust. Aside from the classical matter source $T^{(D)}_{\mu\nu}$, the effective source $\mathcal{T}^{(D)}_{\mu\nu}$ has two distinct components: an ingoing flux (red arrows) with energy of opposite sign to that of the null dust, and the energy distribution deposited in the regular black hole (red shaded region). Both classical matter and effective sources cancel identically at $R=0$. The energy stored in the regular black hole grows while $\dot{M}>0$ (in between the blue dashed lines), decreases while $\dot{M}<0$ (in between the green dashed lines), and remains constant otherwise.
@@ Acknowledgements.
The authors are grateful to Jose Beltrán Jiménez for insightful discussions on 2–dimensional Horndeski gravities, Carlos Barceló, Vitor Cardoso and Stefano Liberati for feedback on the first version of the manuscript, and Ángel Murcia for useful correspondence. RCR acknowledges financial support provided by the Spanish Government through the Ramón y Cajal program (contract RYC2023-045894-I), the Grant No. PID2023-149018NB-C43 funded by MCIN/AEI/10.13039/501100011033, and the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033, as well as the hospitality of the Center of Gravity, a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. VB acknowledges support from the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587, as well as from the Spanish Government through the Grants No. PID2020-118159GB-C43, PID2020-118159GB-C44, PID2023-149018NB-C43 and PID2023-149018NB-C44 (funded by MCIN/AEI/10.13039/501100011033).